Jenna is 4 years older than her sister. the sum of Jenna's age and her sister's age is 32 write an algebraic equation and solve it to find out how old Jenna is
Answer by Theo(13342)

j = jenna's age.
s = her sister's age.

j + s = 32

j = s + 4

replace j with s + 4 to get s + 4 + s = 32

combine like terms to get 2s + 4 = 32

solve for s to get s = (32-4)/2 = 14

her sister's age is 14.

jenna is 4 years older, so jenna is 18.

the sum of their ages is 18 + 14 = 32

solution looks good.

jenna is 18 years old and her sister is 14 years old.

you were asked to find out how old jenna is.

that's 18.